OpenStudy (anonymous):

how do i solve this show work to help me understand 4= 1/2(3d-7)

OpenStudy (anonymous):

Try multiplying both sides by 2 first.

OpenStudy (anonymous):

the 4 and 1/2?

OpenStudy (anonymous):

Yes. Those are both multiplied by 2.

OpenStudy (anonymous):

what about the stuff in the ( ) do i bring them down

OpenStudy (anonymous):

Yep. You haven't done anything with them yet, so the ( ) stuff stays the same. Only when *you* start messing with it does it change.

OpenStudy (anonymous):

ok i have more questions. -10= 2/5(m+4)

OpenStudy (anonymous):

thank you both:)

OpenStudy (anonymous):

Your new question is ambiguous. Which one do you mean: -10 = 2 / (5(m+4)) or -10 = (2/5) * (m+4) ?